# Gravitational wave probes of particle dark matter: a review

### v3

@@ IV.2 Yukawa-like interactions
If DM only couples to gravity, it seems unlikely that GW interferometers could detect this interaction Hall et al. (2018); Du et al. (2023); however, if there is a Yukawa-like coupling between the standard model and DM, allowing for self-interactions as well, DM transiting through the interferometers could very well be detectable. The following potential $V_{i-j}$ is assumed to govern the (self-) interactions between the standard model and DM
@@
[…]
where $\delta_{i,j}$ parametrizes the standard model and DM couplings, $r$ is a length scale, $\lambda$ is the screening length, $s$ is the spin, and $M_{i,j}$ are the two masses. By assuming that the size of DM is much larger than the size of GW detectors and the range of the DM force, but smaller than the average distance between the detectors, the scattering cross section from this process can be derived:
@@
where $r_{\rm DM}$ is the size of the DM objects. While $|\delta_{\rm SM}|<5\times 10^{-4}$ has been strongly constrained by equivalence principle experiments Adelberger et al. (2009), the DM coupling constant could be much larger, as the only constraints come from how DM self-interaction would have influenced structure formation Spergel and Steinhardt (2000) and collisions between galaxies Harvey et al. (2015).
The lack of observed deceleration of DM in the bullet cluster collision constrained the cross-section of DM to be less than $\sim 1$ cm2/g Randall et al. (2008), which provides limits on the coupling constant of self-interacting DM to be:
@@
in which $\log\left[\frac{\lambda}{r_{\text{DM}}}\right]=5$. Based on the current upper limits on $\delta_{\rm SM}$ and $\delta_{\rm DM}$, the coupling of standard-model particles to DM ones could be quite large, and would thus permit an observation of a transit event in GW interferometers. Furthermore, $\delta_{\rm DM}\gg$ 1 would actually alleviate some problems with the cold DM scenario, including overly dense areas of dwarf galaxies found in simulations, and a prediction that the Milky Way produces more stars than are observed Boylan-Kolchin et al. (2011).
IV.3 Projected constraints on Yukawa-like interactions
By using the potential given in Eq. 90, the differential acceleration visible on interferometers can be computed, and from that, the matched-filter signal-to-noise ratio:
[…]
where $a(f)$ is the Fourier transform of the integral of the differential acceleration $a(t)=a_{y}^{\rm EM}-a_{y}^{\rm IM}-(a_{x}^{\rm EM}-a_{x}^{\rm IM})$ between the input mirror (IM) and the end mirror (EM) Saulson (2017), and $S_{n}(f)$ is the acceleration power spectral density of the noise. By using sensitivity curves of LIGO, Virgo and KAGRA and a threshold on $\rho>8$, the number of detections per year of transiting DM with different masses and different screening lengths can be computed, and is shown in Fig. 14 using a power spectral density for advanced LIGO at design sensitivity.
Figure 14: Taken from Hall et al. (2018). Expected number of macroscopic DM transit events per year with $\rho>8$ as measurable with advanced LIGO operating at design sensitivity, as a function of the DM mass, the screening length $\lambda$ and the coupling $g=\delta_{\text{SM}}\delta_{\text{DM}}$ between DM and the standard model.
Further work was done regarding the detection of macroscopic DM in much more detail Du et al. (2023), in which the different kinds of effects of DM on the test masses were derived. Ref. Du et al. (2023) derives the strain on the detector in terms of three observables: (1) the Doppler effect (previously considered in Hall et al. (2018)), (2) the Shapiro delay (light is curved near objects with mass and thus will take a longer time to travel down the interferometer arms, and (3) the Einstein Delay (time runs slower near more massive objects than less massive ones). Unfortunately, all effects are extremely weak because they are solely gravitational, and thus DM would have to couple to the standard model in some way to be detectable. As in Hall et al. (2018), Ref. Du et al. (2023) computes the expected constraints on the effective coupling of DM to the standard model with an effective parameter $\tilde{\alpha}\sim\delta_{i}\delta_{j}$, which are shown in Fig. 15 for a variety of GW detectors and for two choices of screening lengths: $\lambda=1$ m and $\lambda=10^{6}$ m. Existing constraints from the bullet cluster Spergel and Steinhardt (2000), MICROSCOPE Bergé et al. (2018) and neutron-star kinetic heating with and without interactions between DM and baryons that could enhance energy transfer Gresham et al. (2023). While the constraints that could be obtained from GW interferometers do not appear to beat existing ones, it should be noted that if only a part of DM can be explained by a new fifth force, the Bullet cluster constraints on self-interacting DM disappear, while GW bounds fall off linearly with this fraction of DM charged under this fifth force Du et al. (2023).
@@ IV.3 Projected constraints on Yukawa-like interactions
Figure 15: Taken from Du et al. (2023). 90%-projected sensitivity on the standard model/DM interaction parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$ that can be probed from macroscopic DM transiting through several GW experiments, compared to existing constraints from others. One year of observation is assumed, and two choices for the screening length are made. Dotted lines denote existing constraints. Existing constraints are shown in dotted lines. In particular, Bullet cluster and MICROSCOPE constraints are shown in gray Spergel and Steinhardt (2000); Bergé et al. (2018), MICROSCOPE ones. The purple and red lines denote constraints from neutron star kinetic heating Gresham et al. (2023) with and without additional short-range interactions between baryons and DM that would enhance energy transfer.
Given that such deterministic signals may not be detectable in LIGO, Virgo and KAGRA, it was worth considering whether a stochastic background arising from many transiting DM events could be seen Hall et al. (2018); Du et al. (2023). However, again, it is necessary to include the Yukawa interaction to have a chance of detecting a stochastic background.
Finally, another study Lee et al. (2023) considered specifically how heavy transiting DM would appear in KAGRA, as well as the directional dependence of the signal-to-noise ratio, which turns out to be strong due to the geometry of the detector. In KAGRA, it is shown that the suspension system can be particularly sensitive to DM-standard model interactions, between [1,130] Hz.
@@
V GWs from ultralight bosonic clouds around rotating BHs
@@
If a new ultralight boson [^5] In the context of boson clouds around rotating black holes, this new particle could, but need not, be DM exist, it will have a particular mass, frequency and Compton wavelength. Quantum fluctuations could then allow bosons to pop into existence near a rotating BH, whose characteristic length scale is given by half of its Schwarzschild radius. Some bosons will fall in; others will scatter off, and the extent to which either of these scenarios happens depends on the properties of the black hole and the boson. If the following superradiance condition, relating the BH angular frequency $\Omega$ to the boson angular frequency $\omega$ via the magnetic quantum number $m$, is met Brito et al. (2015):
@@
the boson will extract energy from the BH, and thus the “bosonic wave” will be amplified. This amplification will be maximized if the Compton wavelength of this new particle comparable to the characteristic length of a BH. The massive bosons then become bound to the BH, allowing for successive scatterings, thus permitting a huge number of bosons to appear around the BH as a cloud in a given energy state (no limit on the occupation number of these particles exists a priori).. This process is called “superradiance”.
More quantitatively, solving the Klein-Gordon equation for a Kerr BH hole in the presence of a massive scalar field results in a Schrondinger-like equation that exhibits a $1/r$ potential, resulting from the gravitational interaction between the bosons and the BH. Such an equation can be solved to show that the energy states of the boson cloud are quantized analogously to those of the hydrogen atom, allowing us to describe the boson cloud as a “gravitational atom in the sky” Baumann et al. (2019a).
@@
For the black hole to be spun down such that the condition in Eq. 94 is no longer satisfied, scalar and vector clouds require Isi et al. (2019):
@@
where $\chi_{i}$ is the BH spin at birth, and the GW fine-structure constant is
[…]
@@
where $m_{b}$ is the mass of the boson. This new boson could have a spin of 0, 1 or 2, where spin-0 could correspond to the pseudo-scalar QCD axion (or axion-like) particle Brito et al. (2015, 2020), spin-1 could indicate dark photons arising through kinetic mixing with the ordinary photon Siemonsen et al. (2023a), or spin-2 tensor bosons that could arise from modified theories of gravity Brito et al. (2020).
After the cloud builds up, GWs can be emitted in three different ways: (1) through boson-boson annihilation, (2) through boson energy-level transitions, and (3) after a “bosenova”.
@@
It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems would be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki et al. (2015b); Zhu et al. (2020). Furthermore, both galactic BHs Palomba et al. (2019) and remnants of mergers of BHs outside the galaxy could also be detected, the latter of which, in the scalar case, only being visible in ET and Cosmic Explorer Arvanitaki et al. (2010, 2015b); Isi et al. (2019). Annihilating vector boson cloud systems, however, emit stronger, though shorter, GW signals Brito et al. (2015); Baryakhtar et al. (2017); Siemonsen et al. (2023b); Jones et al. (2023, 2025), and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen et al. (2023a).
The boson mass is fixed by the GW frequency, and could in principle take on any ultralight mass. That is why GW experiments across the frequency spectrum have been or are being designed to search for these boson cloud systems. While most of the literature focuses on ground-based GW interferometers, we note that HFGW detectors will also be able to probe BH superradiance of heavier bosons than those detectable by LIGO, Virgo and KAGRA Aggarwal et al. (2021), e.g. Brzeminski et al. (2025), and space-based GW interferometers will probe lighter masses Arvanitaki et al. (2015b); Brito et al. (2017), and space-based detector analyses can be used to aid in discovering boson clouds in the band of ground-based GW interferometers Ng et al. (2020).
As will be discussed, boson cloud self-interactions could have a significant impact on (1) whether the superradiance condition is saturated Yoshino and Kodama (2012, 2015a, 2015b), (2) how quickly a binary black hole system will approach merger Aurrekoetxea et al. (2024a); Takahashi et al. (2024), and (3) the rate of change of the GW frequency of an isolated boson cloud Baryakhtar et al. (2021). Self-interactions would also allow higher energy levels to grow faster. The strength of these interactions will affect the detection prospects in ET and Cosmic Explorer Collaviti et al. (2025), as well as recent search results which have not yet accounted for moderate or strong interactions.

### v4

@@ IV.1 Yukawa-like interactions
If there is a Yukawa-like coupling between the standard model and DM, allowing for self-interactions as well, DM transiting through the interferometers could be detectable. The interaction potential between DM and standard-model particles can be described by the following form:
@@
$\text{where }~i,j=\text{SM,DM.}$~~~~~~~~~~~~
where $\delta_{i,j}$ represents the coupling strengths between particles $i,j$, and $i,j$ could be either standard-model or DM particles. $r$ is a length scale, $\lambda$ is the screening length, $s$ is the spin of the interaction mediator, and $M_{i,j}$ are the two masses. By noting that the physical size of the detector and range of the interaction force are much larger than $r_{\rm DM}$, but much smaller than the separation between DM particles $L_{\rm DM}$, the interaction can be modeled in terms of an effective geometric cross section and derived using the Born approximation Born (1926):
$\frac{d\sigma}{d\Omega}=\frac{M_{\rm DM}^{2}}{4\pi^{2}}\left|\int V(\vec{r})e^{i(\vec{k}_{i}-\vec{k}_{f})\cdot\vec{r}}d^{3}\vec{r}\right|^{2}=\frac{M_{\rm DM}^{2}}{4\pi^{2}}\left|\tilde{V}(\vec{k}_{i}-\vec{k}_{f})\right|^{2}$ (82)
where $\vec{k}_{i}-\vec{k}_{f}$ is the momentum transfer between the initial and final states. The integral is simply the Fourier transform of the potential in Eq. 81 (considering only the interaction term):
$\tilde{V}=\frac{4\pi GM_{\rm DM}^{2}\delta_{\rm DM}^{2}}{(M_{\rm DM}v_{0}\sin({\theta/2}))^{2}+\lambda^{-2}},$ (83)
where $\theta$ is the scattering angle, i.e. the angle between the incoming and outgoing momenta $\vec{k}_{i}$ and $\vec{k}_{f}$, and $\vec{k}_{i}$ and $\vec{k}_{f}$ have been written in terms of $v_{0}$ and $M_{\rm DM}$. Plugging in Eq. 81 into Eq. 82 and integrating over solid angle, the cross-section can be derived:
@@
While $|\delta_{\rm SM}|<5\times 10^{-4}$ has been strongly constrained by equivalence principle experiments Adelberger et al. (2009), the DM coupling constant could be much larger, as the only constraints come from how DM self-interaction would have influenced structure formation Spergel and Steinhardt (2000) and collisions between galaxies Harvey et al. (2015). In particular, the lack of observed deceleration of DM in the bullet cluster collision constrained the cross-section of DM to be less than $\sim 1$ cm2/g Randall et al. (2008), which provides limits on the coupling constant of self-interacting DM to be:
@@
in which $\log\left[\frac{\lambda}{r_{\text{DM}}}\right]=5$, chosen only as a benchmark, corresponding to a force range much larger $r_{\rm DM}$. Based on the current upper limits on $\delta_{\rm SM}$ and $\delta_{\rm DM}$, could allow an observation of a transit event in GW interferometers. Furthermore, $\delta_{\rm DM}\gg$ 1 would actually alleviate some problems with the cold DM scenario, including overly dense areas of dwarf galaxies found in simulations, and a prediction that the Milky Way produces more stars than are observed Boylan-Kolchin et al. (2011).
IV.2 Projected constraints on Yukawa-like interactions
In order to search for transiting macroscopic DM, it has been proposed in Hall et al. (2018) to use matched filtering. To do so, firstly we must calculate the matched filtering SNR $\rho$:
$\rho^{2}=4\int\frac{|a(f)|^{2}}{S_{n}(f)}\,df;$ (86)
Here, $a(f)$ is the Fourier transform of the differential acceleration $a(t)$ between the test masses, and $S_{n}(f)$ is the power spectral density of the detector. The differential acceleration $a(t)$ arises from the DM transiting through the interferometer test masses, and is calculated with the difference in accelerations between the input mirror and end mirror Saulson (2017) in each arm. The form of $a(t)$ can be derived by taking the gradient of Eq. 81:
$\vec{a}(r)=-\frac{GM_{\rm DM}}{r^{2}}\left[1+(-1)^{s}\,\delta_{\rm DM}\,\delta_{\rm SM}\,e^{-r/\lambda}\left(1+\frac{r}{\lambda}\right)\right]\hat{r}$, (87)
and then the individual accelerations along the $x$ and $y$ arms can be computed as:
$\Delta a_{x}$ $=a(x=L_{x},0,0)-a(0,0,0)$ (88)
$\Delta a_{y}$ $=a(0,0=L_{y},0)-a(0,0,0),$ (89)
$\Delta a_{x}$ $=-GM_{\rm DM}\Bigg{[}\frac{L_{x}-x_{\rm DM}}{r_{1}^{3/2}}-\frac{-x_{\rm DM}}{r_{\rm DM}^{3/2}}\Bigg{]}$
$\times\Bigg{[}1+(-1)^{s}\,\delta_{\rm DM}\,\delta_{\rm SM}\,\exp\Big{(}-\frac{r_{1}}{\lambda}\Big{)}\Big{(}1+\frac{r_{1}}{\lambda}\Big{)}\Bigg{]},$ (90)
$\Delta a_{y}$ $=-GM_{\rm DM}\Bigg{[}\frac{L_{y}-y_{\rm DM}}{r_{2}^{3/2}}-\frac{-y_{\rm DM}}{r_{\rm DM}^{3/2}}\Bigg{]}$
$\times\ \Bigg{[}1+(-1)^{s}\,\delta_{\rm DM}\,\delta_{\rm SM}\,\exp\Big{(}-\frac{r_{2}}{\lambda}\Big{)}\Big{(}1+\frac{r_{2}}{\lambda}\Big{)}\Bigg{]},$ (91)
where
$r_{1}$ $=\sqrt{(L_{x}-x_{\rm DM})^{2}+y_{\rm DM}^{2}+z_{\rm DM}^{2}},$ (92)
$r_{2}$ $=\sqrt{x_{\rm DM}^{2}+(L_{y}-y_{\rm DM})^{2}+z_{\rm DM}^{2}},$ (93)
$r_{\rm DM}$ $=\sqrt{x_{\rm DM}^{2}+y_{\rm DM}^{2}+z_{\rm DM}^{2}},$ (94)
and $x_{\rm DM}$, $y_{\rm DM}$ and $z_{\rm DM}$ are the spatial coordinates of the DM object. The acceleration depends directly on the coupling of DM to standard model particles; thus, the SNR will be proportional to this coupling as well. In searches with GW data, a threshold must be set that delineates candidates that are most likely noise with those that could be of astrophysical origin. To estimate the number of detectable transiting DM events, an SNR threshold of $\rho>8$ is applied, which is standard in GW searches Maggiore (2008), and the detection rate as a function of DM mass, and coupling $g\equiv\delta_{\rm SM}\delta_{\rm DM}$ is computed:
$\mathcal{R}\sim\Phi\sigma_{\rm DM-SM}A$ (95)
where $A$ is the effective area of the detector through which DM transits. The effective area for macroscopic DM transits is determined by the physical area of the detector, the angle of incidence of the DM on the detector, and the interaction cross-section between DM and the interferometer. For larger cross sections (larger couplings), the SNR of the signal will increase. Thus, the SNR threshold determines a minimum detectable cross section, which in turns fixes the rate of transiting DM:
We show potential constraints on the rates of DM transiting through advanced LIGO at design sensitivity in Fig. 13. These results demonstrate that macroscopic DM can be detectable, depending on the parameters of the interaction.
Figure 13: Taken from Hall et al. (2018). Expected number of macroscopic DM transit events per year detectable by advanced LIGO at design sensitivity, as a function of DM mass, screening length $\lambda$, and coupling $g=\delta_{\rm SM}\delta_{\rm DM}$. The physics behind this form of DM has been discussed in Section IV.1. The blue, red and green colors indicate different screening lengths $\lambda$ of the Yukawa-like force. The three panels show the rate of transits for DM objects with three different masses. The rate is calculated using Eq. 95 and the preceding equations, assuming a detection threshold on the SNR of 8. The threshold fixes the minimum detectable coupling accessible to a matched filtering search.
Further work in Du et al. (2023) has expanded on this analysis, incorporating additional effects of DM transits, such as the Doppler effect, Shapiro delay, and Einstein delay, each of which contributes to the strain observed in the interferometer. These effects arise from the interaction of the DM object with the interferometer’s test masses, leading to small deviations in the expected strain signal.
  1. 1.
    Doppler effect: As the DM object moves relative to the detector, its velocity induces a periodic shift in the frequency of the signal received by the interferometer. This is similar to the classical Doppler shift seen in sound or light waves, but in this case, it is gravitational. The Doppler effect results in a frequency-dependent modulation of the strain signal, with the amplitude of the strain varying as the velocity of the DM object changes.
  2. 2.
    Shapiro delay: This effect arises due to the curving of spacetime near massive objects. In particular, light traveling near a massive object, such as a transiting DM object, will experience a delay in arrival time. This delay would be observed as a time shift in the signal received by the detectors, as the light takes longer to travel through the curved spacetime near the DM object.
  3. 3.
    Einstein delay: Similar to the Shapiro delay, the Einstein delay occurs due to the effects of gravitational time dilation. Time is slower near more massive objects. Thus, the passage of a massive DM object through the interferometer’s arms would result in a slight delay in the signal, since the clocks at the test masses would tick slower as the object passes by.
While these gravitational effects are real, they are typically very weak unless DM interacts with standard-model particles. To make the effects measurable by GW interferometers, Ref. Du et al. (2023) confirmed that DM must couple with standard-model particles, agreeeing with the findings of Hall et al. (2018). The strength of this coupling is given in Du et al. (2023) by the effective parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$, where $\tilde{\alpha}=g$ in Hall et al. (2018). Thus, detecting these weak gravitational signals requires a sufficiently large value of $\tilde{\alpha}$, which can be constrained by observations of DM transits.
Fig. 14 shows the projected 90% sensitivity on $\tilde{\alpha}$, which can be probed by GW detectors for macroscopic DM. The results, shown for two different screening lengths ($\lambda=1\,\text{m}$ and $10^{6}\,\text{m}$), are compared to existing constraints from the Bullet Cluster Spergel and Steinhardt (2000), MICROSCOPE Bergé et al. (2018), and neutron-star kinetic heating experiments Gresham et al. (2023). While GW detectors may not surpass these existing constraints, they provide complementary information, especially if part of the DM population interacts with the standard model via a new fifth force.
@@ IV.2 Projected constraints on Yukawa-like interactions
Figure 14: Taken from Du et al. (2023). Projected sensitivity at 90% confidence to the standard-model/DM interaction parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$ from macroscopic DM transits through several GW detectors. The physics behind this form of DM has been discussed in Section IV.1. The figure compares the projected sensitivity to existing constraints, including those from the Bullet Cluster, MICROSCOPE, and neutron-star kinetic heating, for two choices of screening length. Here, $T_{\text{obs}}=1$, and $\lambda=1$ m and $\lambda=10^{6}$ m, and an SNR threshold of 2 are assumed. The gray line represents combined limits from Bullet Cluster observations Spergel and Steinhardt (2000); Kahlhoefer et al. (2014) and the MICROSCOPE experiment Bergé et al. (2018); Fayet (2019). Purple and red lines indicate bounds from neutron star kinetic heating Gresham et al. (2023), with and without additional short-range DM–baryon interactions, respectively.
While deterministic signals might be challenging to detect due to the weak gravitational interactions, there is also potential for detecting a stochastic background arising from numerous transiting DM events. This stochastic background could be visible if the Yukawa interaction is sufficiently strong, as explored in Hall et al. (2018); Du et al. (2023). Future work will focus on whether a stochastic signal could be distinguished from the background noise in upcoming GW observations.
@@
V GWs from ultralight boson clouds around rotating black holes
@@
If new ultralight bosons [^8] In the context of boson clouds around rotating black holes, this new particle could, but need not, be DM exist, quantum fluctuations could allow them to pop into existence near a rotating BH. Some bosons will fall in; others will scatter off, and the extent to which either of these scenarios happens depends on the mass and spin of both the black hole and the new boson. If the following so-called “superradiance condition”, which relates the BH angular frequency $\Omega$ to the boson angular frequency $\omega$ via the magnetic quantum number $m$, is met Brito et al. (2015):
@@
the boson will extract energy from the BH, and thus the “bosonic wave” will be amplified. This amplification will be maximized if the Compton wavelength of this new particle is comparable to the radius of a BH. The massive bosons then become bound to the BH because they sit in the potential well of the BH, allowing for successive scatterings, thus permitting a huge number of bosons to appear around the BH as a cloud in a given energy state (no limit on the occupation number of these particles exists). This process is called “superradiance”.
More quantitatively, solving the Klein-Gordon equation for a Kerr BH hole in the presence of a massive scalar field results in a Schrondinger-like equation that exhibits a $1/r$ potential, resulting from the gravitational interaction between the bosons and the BH. The solution of this equation shows that the energy states of the boson cloud are quantized analogously to those of the hydrogen atom, allowing us to describe the boson cloud as a “gravitational atom in the sky” Baumann et al. (2019a).
@@
The cloud will continue to build up until the BH is spun down enough such that the condition in Eq. 96 is no longer satisfied. In particular, the timescales $\tau_{\rm inst}^{\rm(s)}$ and $\tau_{\rm inst}^{\rm(v)}$ for which this occurs scalar and vector clouds, respectively, are Isi et al. (2019) [^9] Note that we do not consider tensor boson clouds in this work:
@@
where $\chi_{i}$ is the BH spin at birth, and the GW fine-structure constant $\alpha$ is
$\alpha$ $=GM_{\rm BH}m_{b}$ (99)
$=0.11\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{m_{b}}{1.47\times 10^{-12}\,\mathrm{eV}}\right)$ (100)
where $m_{b}$ is the mass of the boson and $M_{\rm BH}$ is the mass of the BH.
@@
This new boson could have a spin of 0, 1 or 2, where spin-0 could correspond to the pseudo-scalar QCD axion (or axion-like) particle Brito et al. (2015, 2020), spin-1 could indicate dark photons arising through kinetic mixing with the ordinary photon Siemonsen et al. (2023a) [^10] Note that this dark photon is not the dark photon described in Section II.4., or spin-2 tensor bosons that could arise from modified theories of gravity Brito et al. (2020). After the cloud builds up in an isolated BH system, GWs can be emitted in three different ways: (1) through boson-boson annihilation, (2) through boson energy-level transitions, and (3) after a “bosenova”, all of which will be discussed in Sections V.1, V.2 and V.4. Furthermore, the GW frequency is fixed by the boson mass, which could in principle take on any ultralight mass; thus, it is important to note that GW interferometers, from pulsar timing arrays to HFGW instruments, will allow an extensive mass range to be probed, from $10^{-22},10^{-9}]\,\mathrm{eV}$ Aggarwal et al. (2021); Brzeminski et al. (2025); Arvanitaki et al. (2015a); Brito et al. (2017); Ng et al. (2020).
@@
It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems could be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki et al. (2015a); Zhu et al. (2020); and if such sources are not seen, particular boson masses can be disfavored – this will be the subject of Section VI. Furthermore, both galactic BHs Palomba et al. (2019) and remnants of mergers of BHs outside the galaxy could also be detected, the latter of which, in the scalar case, only being visible in ET and Cosmic Explorer Arvanitaki et al. (2010, 2015a); Isi et al. (2019), which will be discussed more in Section VI.1. Annihilating vector boson cloud systems, however, emit stronger, though shorter, GW signals Brito et al. (2015); Baryakhtar et al. (2017); Siemonsen et al. (2023b); Jones et al. (2023, 2025), and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen et al. (2023a).
In this section, we will discuss the three main ways in which GWs can be emitted from isolated BH/ boson cloud systems: annihilation, transitions and bosenovae (Sections V.1, V.2 and V.4). We will also detail how self-interactions between the bosons could impact the superradiance process and the resulting GWs in Section V.3. Finally, we will explain some of the rich physics that can occur when boson clouds form around binary black hole systems in Section V.5.